A force 25N acts on a body of mass 6kg, the direction of which forms the angle 60⁰ with the horizontal direction, a. Find the acceleration with which the body moves. Friction is not taken into account. b. Find the acceleration when the coefficient of friction = 0.1


1
Expert's answer
2021-02-09T10:42:10-0500

a)


"F_x=ma_x\to a_x=F_x/m=25\cdot\cos60°/6=2.083\ (m/s^2)"


b)


"F_x-\mu N=ma_x\to a_x=(F_x-\mu N)/m"


"N=mg-25\cdot\sin60°=6\cdot9.8-25\cdot\sin60°=37.15(N)"


"a_x=(F_x-\mu N)/m=(25\cdot \cos60°-0.1\cdot 37.15)/6=1.46\ (m/s^2)"